OpenStudy (anonymous):

Use the substitution method to solve the following system of equations 8x + y = 10 6x + 5y = –1

OpenStudy (anonymous):

The substitution method involves solving for a variable in terms of the other one. 8x+y=10 y=10-8x Now you'd substitute this y-value into the second equation. 6x+5(10-8x)=-1 Do you follow?

OpenStudy (abb0t):

I don't really follow. Can you elaborate more, please.

OpenStudy (anonymous):

Next you'd solve for the x-value. 6x+50-40x=-1 -34x+50=-1 -34x=-51 x=-51/-34 x=1.5 Now you'd plug in this value into either equation (for x) and solve for y.